William Howe must decide whether to start a business renting beach umbrellas at an ocean resort during June, July and August of next summer. He believes he can rent each umbrella to vacationers at $5 a day, and he intends to lease 50 umbrellas for the three- month period for$3,000. To operate this business, he does not have to hire anyone (but himself), and he has no expenses other than the leasing costs and a fee of $3,000 per month to rent the business location. Howe is a college student, and if he did not operate this business, he could earn $4,000 for the Three- month period doing construction work.


1
Expert's answer
2021-04-28T20:15:13-0400

Total problem:

William Howe must decide whether to start a business renting beach umbrellas at an ocean resort during June, July, and August of next summer. He believes he can rent each umbrella to vacationers at $5 a day, and he intends to lease 50 umbrellas for the three- month period for $3,000. To operate this business, he does not have to hire anyone (but himself), and he has no expenses other than the leasing costs and a fee of $3,000 per month to rent the business location. Howe is a college student, and if he did not operate this business, he could earn $4,000 for the three-month period doing construction work. If there are 80 days during the summer when beach umbrellas are demanded and Howe rents all 50 of his umbrellas on each of these days, what will be his accounting profit t for the summer?

Answer:

Accounting profit = Revenue – Expenses

Revenue = Number of days "\times" Number of umbrellas "\times" Rent per day for an umbrella

Revenue "= 80 \times 50 \times 5 = 20000"

Expenses = Lease cost + Rent cost

Expenses "= 3000 + (3 \times 3000) = 12000"

Accounting profit = 20000 – 12000 = 8000